# FiCo-ITR: bridging fine-grained and coarse-grained image-text retrieval for comparative performance analysis

### v1

@@ 1 Introduction
Cross-Modal Retrieval (CMR) involves using one type of data, such as text, to search for another type of data, such as images. Unlike general multi-modal tasks, CMR specifically focuses on bridging the gap between different modalities to enable retrievals across modalities. CMR has gained prominence over the past decade due to its success in various applications, including e-commerce [1, 2], content-based retrieval [3, 4], video surveillance [5], and recommendation systems [6]. When the retrieval task specifically involves images and text, it is referred to as Image-Text Retrieval (ITR). There are two distinct subfields within ITR: Fine-Grained (FG) and Coarse-Grained (CG) ITR.
@@
FG ITR aims to find instance-level matches, retrieving the image that directly corresponds to a detailed text query, and vice versa. State-of-the-art FG methods employ large-scale Vision-Language Pretraining (VLP) followed by retrieval fine-tuning, often through the use of contrastive learning techniques [7].
@@
CG ITR focuses on category-level retrieval, where the retrieved samples should broadly belong to the semantic category which the query is searching for, rather than aiming for specific exact matches. State-of-the-art CG methods implement Cross-Modal Hashing [8], which train hash functions to map image and text samples onto a common Hamming subspace for efficient bitwise similarity comparisons. Figure 1 illustrates the different search criteria between FG and CG search approaches.
@@
Figure 1: Comparison of Fine-Grained (FG) and Coarse-Grained (CG) Image-Text Retrieval approaches. FG search uses continuous embeddings, aiming to find the retrieval sample that specifically corresponds to the query sample. Under evaluation conditions, this involves finding the retrieval sample with the same ID as the query sample. CG search employs bitwise hash codes to find retrieval samples that are more broadly relevant to the query instead of exact matches. During evaluation, this involves finding any retrieval sample with at least one matching category label relative to the query’s relevant category labels. The broader search criteria of CG search allows for more efficient computational costs, as seen in the comparison of encoding time and embedding storage costs of two representative FG and CG models (IMRAM [9] and UCCH [10])
@@
Despite sharing the same overarching retrieval task, the subfields of FG and CG ITR have evolved independently with limited integration, leading to the following three major challenges: 1) Because of methodological differences between the two subfields, obtaining informative quantitative results which are directly comparable is a non-trivial task. 2) Researchers in one subfield rarely benchmark against works from the other in a comprehensive manner. Although FG and CG methods have been comparatively surveyed within the literature [11], direct empirical comparative evaluations of recent representative models are lacking. 3) Traditional ITR benchmark datasets such as Flickr30K [12] (1K sample test set) and MS-COCO [13] (5K sample test set) are small compared to real-world applications. Therefore, these benchmark datasets do not offer a comprehensive understanding of model performance on large-scale data, potentially skewing perceived retrieval performance-efficiency trade-offs. Obtaining insights into the comparative performance and tradeoffs of FG and CG methods is crucial for informing the use cases of each method and exploring the potential to leverage their strengths by integrating both types of methods into a hybrid approach. The contributions of this paper directly address the three identified challenges as follows:
@@
    A FiCo-ITR library and unified toolkit that standardises evaluation methodologies for the otherwise disjoint CG and FG ITR subfields. The implementation addresses the challenge of establishing a fair comparative framework by carefully selecting tasks, datasets, and metrics which appropriately address the distinct characteristics of each subfield.
@@
    An empirical evaluation of the recall, precision, encoding time, storage costs, query-time attention costs and similarity search time of representative CG and FG approaches, identifying the relative strengths and limitations of the two subfields using the proposed library.
@@
    Scalability experiments using increasing amounts of data that provide insights into the real-world scalability of the selected models. These experiments offer a perspective on the balance between retrieval performance and scalability which is often overlooked when evaluating on default traditional, relatively small benchmark datasets.
@@ 2.1 Fine-grained image-text retrieval
FG methods aim to map visual and textual information to a joint space where relevant samples are aligned at the instance level. The alignment of samples is achieved through learning shared feature representations from image-text pairs during training. Establishing the joint space enables retrieval across modalities, where the relative distance of items in the shared space determines their relevance. While FG methods share the same fundamental principles as CG ones, they differ in the level of scrutiny with which the processed samples are analysed. Specifically, FG methods focus on low-level features and object-level relationships to achieve an in-depth understanding of the scene within a given sample. The format of the encoded samples is typically real-valued continuous embeddings. These embeddings enable calculating the similarity between samples using vector distance measurements such as Cosine distance [14].
@@
Traditional fine-grained methods. Early deep learning-based FG methods typically employed Long Short-Term Memory (LSTM) networks for text encoding and convolutional neural networks (CNNs) such as AlexNet and VGGNet for image feature extraction [15]. VSE++ [14] built on this approach by experimenting with VGG19 [16] and ResNet152 [17] for image encoding along with a GRU-based text encoder and implementing hard negative mining and reranking loss functions. SCAN [18] achieved a breakthrough in performance through its proposed bottom-up attention module aligning image regions and words based on Faster R-CNN object detection [19]. CAMP [20] aggregates salient messages between image regions and words via attention to handle negative pairs before directly predicting matching scores. VSRN [21] introduced hierarchical reasoning to capture region-concept interactions and was later upgraded to VSRN++ [22] using BERT [23] text features. KASCE [24] expands image concepts using “common-sense” relationships from scene graphs and selects the most relevant expansions. CAAN [25] employs context-aware attention to selectively attend to fragments based on inter- and intra-modal semantics. IMRAM [9] progressively aligns fragments through a recurrent attention unit to capture different semantics at each step, alongside a memory unit to accumulate cues.
@@
Fine-grained vision-language pretraining methods. Following the success of the attention mechanism in BERT [23], the transformer architecture has been extensively adapted for vision-language tasks. This adaptation predominantly involves leveraging pretraining on large-scale data to obtain task-agnostic features for subsequent fine-tuning onto downstream tasks such as ITR. Initial VLP methods are categorised into two main encoder architectures: Dual-encoder and fusion-encoder architectures. Dual-encoders, such as ALIGN [26] and BEIT-3 [27], employ separate image and text encoders, allowing the independent computation and offline storage of embeddings for each modality. This approach eliminates the need to compute embeddings at query time. Fusion-encoder models, such as UNITER [28] and METER [29], enhance interactivity between modalities by employing a unified encoder that jointly processes image and text inputs. Although the fusion encoder approach may potentially lead to higher-quality embeddings, the embeddings must be computed at query time due to the requirement to process all possible query/retrieval sample combinations jointly. Building on these two approaches, models such as X2-VLM [30] and BLIP-2 [31] adopt a hybrid methodology. These hybrid methods first leverage a dual-encoder step to independently compute embeddings to boost efficiency. Then, to further improve the embedding quality, they employ a fusion-encoder reranking step that allows for modality interaction. Specifically, the top candidates retrieved from the initial dual-encoder step are reranked using the fusion-encoder during query time.
@@
Efficiency in FG ITR. To address efficiency challenges inherent to FG ITR, recent approaches have aimed to optimise inference speeds by proposing lightweight implementations of the dual-encoder architecture. LightningDOT [32] tackles this by simplifying pre-training tasks, maximising the amount of computations that can be done offline, and promoting the use of lightweight encoders. VLDeformer [33] proposes a two-stage retrieval process, first using a transformer learning stage followed by more efficient indexing-based retrieval in the second stage. HiVLP [34] uses coarse-grained screening as a first step, followed by a fine-grained rerank in the second step. However, rather than a hash-based approach, HiVLP generates features for the coarse step using early layers of the transformer stack, whereas fully inferred features from later layers are used for the fine-grained rerank.
@@ 2.2 Coarse-grained image-text retrieval
CG methods, similarly to FG methods, also aim to learn joint visual and textual representations. However, in contrast to the instance-level alignment of FG methods, CG methods aim to align relevant samples at a broader semantic category level. By using a broader criterion, CG methods can place more emphasis on computational efficiency. The most prominent approach within CG ITR is Cross-Modal Hashing (CMH) [8], which is primarily characterised by the use of bit hash codes to represent their encoded data. The use of hash codes aims for lower storage costs and faster retrieval speeds, due to the bit hash format being inherently lightweight and the associated bitwise operations being less computationally complex than continuous embedding operations. Recent CMH methods use deep neural networks to learn hash functions that map image and text samples to binary hash codes. During retrieval, the hash codes are compared using Hamming distance, an efficient similarity measure for bit strings which counts the number of differing bits between two equal-length bit hash codes.
@@
Supervised cross-modal hashing. Supervised CMH methods leverage multi-category labelling to train the hash function for each modality. DCMH [35] pioneered this approach by proposing an end-to-end deep learning CMH framework. Leveraging Generative Adversarial Networks (GAN), SSAH [36] implements label information as network input to strengthen category alignment in the hash space. AGAH [37] uses label information directly in its loss function, implementing a multi-labeling map. DADH [38] adopts a weighted cosine triplet-margin constraint for ranking-based similarity preservation. DCHUC [39] introduces a four-step iterative optimisation process that allows simultaneous learning of unified hash codes for database samples and modality-specific hashing functions for unseen queries. DCHMT [40] employs a dual transformer tower network and a differentiable hashing module, enabling location-aware image encoding and continuous, gradient descent-optimised modality representation. LDSAH [41] integrates label-wise semantic alignment with a dissimilarity-penalising strategy using a combination of Jensen–Shannon divergence loss and attention-driven sample re-weighting.
@@
Unsupervised cross-modal hashing. Unsupervised CMH methods use image-text pair coupling information to learn the modality hash functions, avoiding the reliance on category labelling information. This makes unsupervised CMH methods more analogous to typical FG methods, which also do not rely on labelling. Unsupervised UGACH [42] uses GANs to exploit the underlying manifold structure of cross-modal data with a max-margin ranking loss. UDCMH [43] constructs self-taught deep hash functions by minimising quantisation errors while preserving nearest and farthest neighbourhood relationships. DJSRH [44] uses a joint semantics affinity matrix to combine neighbourhood relations from different modalities for improved quantisation and batch-wise training efficiency. DSAH [45] employs a semantic-alignment loss function with auto-encoder-based hash code generation. JDSH [46] proposes a similarity decision and weighting approach which uses a threshold-based weighting scheme to increase the discrimination of hash codes. DGCPN [39] uses a graph-based framework with nodes representing related data and employs a conditional probability model to evaluate the coherence between neighbouring nodes. ADV [47] adopts fine-grained learning objectives for a two-step instance-level retrieval task where shorter hash code embeddings perform initial screening for subsequent reranking using longer hash codes. UCCH [10] uses contrastive learning with a momentum optimiser and introduces a cross-modal ranking learning loss to address binary-continuous relaxation challenges and mitigate the impact of false-negative pairs.
@@ 3 Proposed library and toolkit
Figure 2: An extendable framework of the FiCo-ITR library and toolkit. The pipeline consists of five main components: 1) Data Pre-Processing, which standardises dataset handling and offers optional label generation via Query2Label (Q2L) [48] for unlabeled datasets; 2) Model Encoding, supporting embeddings in the form of binary hash codes and continuous embeddings, as well as direct similarity matrices; 3) Similarity Measures, implementing four distance measures for uniform similarity calculation; 4) Retrieval Tasks, implementing instance- and category-level retrieval for both (i $\rightarrow$ t) and (t $\rightarrow$ i); and 5) Evaluation Metrics, reporting Recall@K for instance-level retrieval and mAP@K with P/R curves for category-level retrieval
@@
1) Data pre-processing. To enable both instance-level and category-level retrieval evaluations, the selected benchmark datasets must have full-sentence captions as well as semantic category labelling. For datasets which do not have semantic category labelling, we employ the Query2Label (Q2L) [48] classifier to label them with synthetic semantic category labels. We adopt pre-processing steps which most closely adhere to each subfield’s standard approaches [10, 18]: For CG models, FiCo-ITR provides the toolkit to extract 4096D VGG-19 features for images and 300D Doc2Vec features for text. For traditional FG models, we refer to the Bottom-Up Attention [49] repository for Faster RCNN 1024x36D features for images, whereas the raw text is given as input. VLP models process raw data using their respective transformer encoder stacks.
2) Model encoding. Given the pre-processed data as input, the samples chosen for each split must be consistent across all evaluated models to ensure the most directly comparable evaluation. To this end, the FiCo-ITR toolkit provides functionality to split samples into pre-defined splits along their IDs to ensure consistency of results when comparing different models. In the case of extending the use of FiCo-ITR to other datasets not included in this study, FiCo-ITR provides the tools to stratify, record ID designation, and align splits across models. Three embedding output types are handled: Binary hash codes, continuous embeddings, or in the case of models which employ query-time attention, directly computed similarity matrices.
@@
3) Similarity measures. Given a set of evaluation embeddings, we implement four similarity measures for similarity rank ordering: For traditional binary hash codes (0s and 1s), we implement Hamming distance which calculates the number of differing bits between two hash codes efficiently on CPUs. However, the library also supports hash codes represented as -1s and 1s, enabling the use of inner product, cosine similarity, and Euclidean distance. This design choice allows users to leverage GPU acceleration for hash code comparisons, potentially boosting performance in GPU-centric environments. For continuous embeddings, all measures except Hamming distance are applicable. The consistency in similarity computation within this step ensures the output similarity matrices are aligned for the subsequent tasks.
@@
$R@k=\frac{\text{number of relevant items in top }k\text{ results}}{\text{totalrelevant items}}$ (1)
Recall is particularly suitable for instance-level retrieval as it measures whether specific, individual items have been found. For category-level retrieval, we use mean average precision at $k$ (mAP@$k$) as the standard evaluation metric. Precision at $k$ ($P@k$) is defined as follows:
$P@k=\frac{\text{number of relevant items in top }k\text{ results}}{\text{number of retrieved items}}$ (2)
For a single query, AP@$k$ averages the precision values at all available ranks within the top $k$. For a query set, mAP@$k$ computes the mean of the AP@$k$ scores of all the queries within the query set. mAP@$k$ is well-suited for evaluating category-level retrieval where queries typically have a large number of relevant items, as it measures a model’s ability to include many relevant results towards the top of a large retrieved set. Additionally, we implement the 11-point precision-recall curve for the category-level retrieval task, which plots interpolated precision at recall levels $R=\{0.0,0.1,\ldots,1.0\}$.

### v2

@@ 1 Introduction
Cross-Modal Retrieval (CMR) involves using one type of data, such as text, to search for another type of data, such as images. Unlike general multi-modal tasks, CMR specifically focuses on bridging the gap between different modalities to enable retrievals across modalities. CMR has gained prominence over the past decade due to its success in various applications, including e-commerce [Ma2022EICLIPEI, Zhang2021FashionFM], content-based retrieval [nakatsuka2023content, gong2023neural], video surveillance [yang2024detection], and recommendation systems [Truong2021ExploringCU]. When the retrieval task specifically involves images and text, it is referred to as Image-Text Retrieval (ITR). There are two distinct subfields within ITR: Fine-Grained (FG) and Coarse-Grained (CG) ITR.
@@
FG ITR aims to find instance-level matches, retrieving the image that directly corresponds to a detailed text query, and vice versa. State-of-the-art FG methods employ large-scale Vision-Language Pretraining (VLP) followed by retrieval fine-tuning, often through the use of contrastive learning techniques [gan2022vision].
@@
CG ITR focuses on category-level retrieval, where the retrieved samples should broadly belong to the semantic category which the query is searching for, rather than aiming for specific exact matches. State-of-the-art CG methods implement Cross-Modal Hashing [luo2023survey], which train hash functions to map image and text samples onto a common Hamming subspace for efficient bitwise similarity comparisons. Figure 1 illustrates the different search criteria between FG and CG search approaches.
@@
Figure 1: Comparison of Fine-Grained (FG) and Coarse-Grained (CG) Image-Text Retrieval approaches. FG search uses continuous embeddings, aiming to find the retrieval sample that specifically corresponds to the query sample. Under evaluation conditions, this involves finding the retrieval sample with the same ID as the query sample. CG search employs bitwise hash codes to find retrieval samples that are more broadly relevant to the query instead of exact matches. During evaluation, this involves finding any retrieval sample with at least one matching category label relative to the query’s relevant category labels. The broader search criteria of CG search allows for more efficient computational costs, as seen in the comparison of encoding time and embedding storage costs of two representative FG and CG models (IMRAM [chen2020imram] and UCCH [hu2022unsupervised])
@@
Despite sharing the same overarching retrieval task, the subfields of FG and CG ITR have evolved independently with limited integration, leading to the following three major challenges: 1) Because of methodological differences between the two subfields, obtaining informative quantitative results which are directly comparable is a non-trivial task. 2) Researchers in one subfield rarely benchmark against works from the other in a comprehensive manner. Although FG and CG methods have been comparatively surveyed within the literature [zhu2023cross], direct empirical comparative evaluations of recent representative models are lacking. 3) Traditional ITR benchmark datasets such as Flickr30K [plummer2015flickr30k] (1K sample test set) and MS-COCO [lin2014microsoft] (5K sample test set) are small compared to real-world applications. Therefore, these benchmark datasets do not offer a comprehensive understanding of model performance on large-scale data, potentially skewing perceived retrieval performance-efficiency trade-offs.
The lack of standardised evaluation methodologies and empirical comparisons between FG and CG methods has hindered the field’s understanding of their relative strengths and limitations in real-world scenarios. This understanding is necessary for informing model selection decisions and identifying opportunities for hybrid approaches that could leverage the advantages of both methodologies. To address these critical gaps, this paper makes the following contributions:
@@
    The implementation and distribution of the FiCo-ITR library and unified toolkit that bridges the methodological differences between the CG and FG ITR subfields. This contribution addresses the challenge of establishing fair comparisons by implementing carefully selected tasks, datasets, and metrics that accommodate the distinct characteristics of each subfield while maintaining evaluation consistency. The library is made available in the Python Package Index (PyPI) and GitHub, enabling researchers to conduct standardised evaluations across both subfields, facilitating more informed comparisons and potential cross-pollination of ideas between the two approaches.
@@
    A systematic empirical evaluation that quantifies the trade-offs between CG and FG approaches across multiple aspects of performance: Recall, precision, encoding time, storage costs, query-time attention costs and similarity search time. This evaluation provides the first comprehensive empirical basis for understanding the relative strengths and limitations of representative approaches from both subfields, addressing the current lack of direct comparative analysis in the literature.
@@
    Scalability experiments that evaluate model performance using incrementally larger retrieval sets, quantifying the computational costs of large-scale ITR. These experiments reveal critical trade-offs between retrieval performance and computational efficiency that only become apparent at scale, providing practical guidance for real-world deployment decisions.
@@ 2.1 Fine-grained image-text retrieval
FG methods aim to map visual and textual information to a joint space where relevant samples are aligned at the instance level. The alignment of samples is achieved through learning shared feature representations from image-text pairs during training. Establishing the joint space enables retrieval across modalities, where the relative distance of items in the shared space determines their relevance. While FG methods share the same fundamental principles as CG ones, they differ in the level of scrutiny with which the processed samples are analysed. Specifically, FG methods focus on low-level features and object-level relationships to achieve an in-depth understanding of the scene within a given sample. The format of the encoded samples is typically real-valued continuous embeddings. These embeddings enable calculating the similarity between samples using vector distance measurements such as Cosine distance [faghri2017vse++].
@@
Traditional fine-grained methods. Early deep learning-based FG methods typically employed Long Short-Term Memory (LSTM) networks for text encoding and convolutional neural networks (CNNs) such as AlexNet and VGGNet for image feature extraction [kiros2014unifying]. VSE++ [faghri2017vse++] built on this approach by experimenting with VGG19 [simonyan2014very] and ResNet152 [he2016deep] for image encoding along with a GRU-based text encoder and implementing hard negative mining and reranking loss functions. SCAN [lee2018stacked] achieved a breakthrough in performance through its proposed bottom-up attention module aligning image regions and words based on Faster R-CNN object detection [ren2015faster]. CAMP [wang2019camp] aggregates salient messages between image regions and words via attention to handle negative pairs before directly predicting matching scores. VSRN [li2019visual] introduced hierarchical reasoning to capture region-concept interactions and was later upgraded to VSRN++ [li2022image] using BERT [devlin2019bert] text features. KASCE [shi2019knowledge] expands image concepts using “common-sense” relationships from scene graphs and selects the most relevant expansions. CAAN [zhang2020context] employs context-aware attention to selectively attend to fragments based on inter- and intra-modal semantics. IMRAM [chen2020imram] progressively aligns fragments through a recurrent attention unit to capture different semantics at each step, alongside a memory unit to accumulate cues.
@@
Fine-grained vision-language pretraining methods. Following the success of the attention mechanism in BERT [devlin2019bert], the transformer architecture has been extensively adapted for vision-language tasks. Moreover, first pioneered by the model ViLT [kim2021vilt], state-of-the-art performance is achieved by using transformer encoder stacks end-to-end, without the need for additional pre-processing steps such as object detection feature extraction predominant in traditional FG models. This shift most notably involves leveraging pretraining on large-scale data to obtain task-agnostic features for subsequent fine-tuning onto downstream tasks such as ITR. Initial VLP methods are categorised into two main encoder architectures: Dual-encoder and fusion-encoder architectures. Dual-encoders, such as ALIGN [jia2021scaling] and BEIT-3 [wang2023image], employ separate image and text encoders, allowing the independent computation and offline storage of embeddings for each modality. This approach eliminates the need to compute embeddings at query time. Fusion-encoder models, such as UNITER [chen2020uniter] and METER [dou2022empirical], enhance interactivity between modalities by employing a unified encoder that jointly processes image and text inputs. Although the fusion encoder approach may potentially lead to higher-quality embeddings, the embeddings must be computed at query time due to the requirement to process all possible query/retrieval sample combinations jointly. Building on these two approaches, models such as X2-VLM [zeng2023x] and BLIP-2 [li2023blip] adopt a hybrid methodology. These hybrid methods first leverage a dual-encoder step to independently compute embeddings to boost efficiency. Then, to further improve the embedding quality, they employ a fusion-encoder reranking step that allows for modality interaction. Specifically, the top candidates retrieved from the initial dual-encoder step are reranked using the fusion-encoder during query time.
@@
Efficiency in FG ITR. To address efficiency challenges inherent to FG ITR, recent approaches have aimed to optimise inference speeds by proposing lightweight implementations of the dual-encoder architecture. LightningDOT [sun2021lightningdot] tackles this by simplifying pre-training tasks, maximising the amount of computations that can be done offline, and promoting the use of lightweight encoders. VLDeformer [zhang2022vldeformer] proposes a two-stage retrieval process, first using a transformer learning stage followed by more efficient indexing-based retrieval in the second stage. HiVLP [chen2022hivlp] uses coarse-grained screening as a first step, followed by a fine-grained rerank in the second step. However, rather than a hash-based approach, HiVLP generates features for the coarse step using early layers of the transformer stack, whereas fully inferred features from later layers are used for the fine-grained rerank.
@@ 2.2 Coarse-grained image-text retrieval
CG methods, similarly to FG methods, also aim to learn joint visual and textual representations. However, in contrast to the instance-level alignment of FG methods, CG methods aim to align relevant samples at a broader semantic category level. By using a broader criterion, CG methods can place more emphasis on computational efficiency. The most prominent approach within CG ITR is Cross-Modal Hashing (CMH) [luo2023survey], which is primarily characterised by the use of bit hash codes to represent their encoded data. The use of hash codes aims for lower storage costs and faster retrieval speeds, due to the bit hash format being inherently lightweight and the associated bitwise operations being less computationally complex than continuous embedding operations. Recent CMH methods use deep neural networks to learn hash functions that map image and text samples to binary hash codes. During retrieval, the hash codes are compared using Hamming distance, an efficient similarity measure for bit strings which counts the number of differing bits between two equal-length bit hash codes.
@@
Supervised cross-modal hashing. Supervised CMH methods leverage multi-category labelling to train the hash function for each modality. DCMH [jiang2017deep] pioneered this approach by proposing an end-to-end deep learning CMH framework. Leveraging Generative Adversarial Networks (GAN), SSAH [li2018self] implements label information as network input to strengthen category alignment in the hash space. AGAH [gu2019adversary] uses label information directly in its loss function, implementing a multi-labeling map. DADH [bai2020deep] adopts a weighted cosine triplet-margin constraint for ranking-based similarity preservation. DCHUC [tu2020deep] introduces a four-step iterative optimisation process that allows simultaneous learning of unified hash codes for database samples and modality-specific hashing functions for unseen queries. DCHMT [tu2022differentiable] employs a dual transformer tower network and a differentiable hashing module, enabling location-aware image encoding and continuous, gradient descent-optimised modality representation. LDSAH [li2023label] integrates label-wise semantic alignment with a dissimilarity-penalising strategy using a combination of Jensen–Shannon divergence loss and attention-driven sample re-weighting.
@@
Unsupervised cross-modal hashing. Unsupervised CMH methods use image-text pair coupling information to learn the modality hash functions, avoiding the reliance on category labelling information. This makes unsupervised CMH methods more analogous to typical FG methods, which also do not rely on labelling. Unsupervised UGACH [zhang2018unsupervised] uses GANs to exploit the underlying manifold structure of cross-modal data with a max-margin ranking loss. UDCMH [wu2018unsupervised] constructs self-taught deep hash functions by minimising quantisation errors while preserving nearest and farthest neighbourhood relationships. DJSRH [su2019deep] uses a joint semantics affinity matrix to combine neighbourhood relations from different modalities for improved quantisation and batch-wise training efficiency. DSAH [yang2020deep] employs a semantic-alignment loss function with auto-encoder-based hash code generation. JDSH [liu2020joint] proposes a similarity decision and weighting approach which uses a threshold-based weighting scheme to increase the discrimination of hash codes. DGCPN [tu2020deep] uses a graph-based framework with nodes representing related data and employs a conditional probability model to evaluate the coherence between neighbouring nodes. ADV [li2021efficient] adopts fine-grained learning objectives for a two-step instance-level retrieval task where shorter hash code embeddings perform initial screening for subsequent reranking using longer hash codes. UCCH [hu2022unsupervised] uses contrastive learning with a momentum optimiser and introduces a cross-modal ranking learning loss to address binary-continuous relaxation challenges and mitigate the impact of false-negative pairs.
@@ 3 Proposed library and toolkit
Figure 2: An extendable framework of the FiCo-ITR library and toolkit. The pipeline consists of five main components: 1) Data Pre-Processing, which standardises dataset handling and offers optional label generation via Query2Label (Q2L) [liu2021query2label] for unlabeled datasets; 2) Model Encoding, supporting embeddings in the form of binary hash codes and continuous embeddings, as well as direct similarity matrices; 3) Similarity Measures, implementing four distance measures for uniform similarity calculation; 4) Retrieval Tasks, implementing instance- and category-level retrieval for both (i $\rightarrow$ t) and (t $\rightarrow$ i); and 5) Evaluation Metrics, reporting Recall@K for instance-level retrieval and mAP@K with P/R curves for category-level retrieval
@@
1) Data pre-processing. To enable both instance-level and category-level retrieval evaluations, the selected benchmark datasets must have full-sentence captions as well as semantic category labelling. For datasets which do not have semantic category labelling, we employ the Query2Label (Q2L) [liu2021query2label] classifier to label them with synthetic semantic category labels. We adopt pre-processing steps which most closely adhere to each subfield’s standard approaches [hu2022unsupervised, lee2018stacked]: For CG models, FiCo-ITR provides the toolkit to extract 4096D VGG-19 features for images and 300D Doc2Vec features for text. Images are preprocessed by resizing to VGG-19’s expected 224×224 pixels and applying standard normalisation (mean=[0.485, 0.456, 0.406], std=[0.229, 0.224, 0.225]). The VGG-19 network, pretrained on ImageNet, has its final three layers removed to obtain the feature vectors. Text features are generated using Doc2Vec trained on the combined training and test captions, with minimum word count thresholding to handle vocabulary sparsity. For traditional FG models, we refer to the Bottom-Up Attention [Anderson2017up-down] repository for Faster R-CNN to extract 36 region proposals per image with 1024-dimensional features, whereas the raw text is given as input to each model’s specific encoder. VLP models process raw data using their respective transformer encoder stacks.
2) Model encoding interface. The framework provides standardised interfaces for handling model outputs while maintaining evaluation consistency across different encoding approaches. While the actual encoding process is model-specific and outside the scope of our framework, FiCo-ITR implements pre-encoding dataset management and post-encoding output handling. For dataset management, the toolkit provides functionality to split samples into pre-defined splits along their IDs, ensuring that when different models encode the same test set, they process exactly the same samples in the same order. For output handling, the framework supports three types of encoded outputs: Binary hash codes from CG models (typically 64-bit to 2048-bit lengths), continuous embeddings from FG models (typically 256 to 2048 dimensions), and directly computed similarity matrices from models employing query-time attention. This flexible interface design enables fair comparisons while respecting each model’s native encoding process.
3) Similarity measures. Given a set of evaluation embeddings, we implement four similarity measures optimised for different embedding types:
  • Cosine similarity for continuous embeddings, computed as the normalised dot product between vectors $\mathbf{x}$ and $\mathbf{y}$: $\frac{\mathbf{x}\cdot\mathbf{y}}{\|\mathbf{x}\|\|\mathbf{y}\|}$, with values ranging from [-1, 1]
  • Euclidean similarity, transformed to a (0, 1] range through $1/(1+d+\epsilon)$ where $d$ is the Euclidean distance $\sqrt{\sum(\mathbf{x}-\mathbf{y})^{2}}$ and $\epsilon=10^{-8}$ for numerical stability
  • Hamming similarity for binary hash codes, computed as the negative normalised count of differing bits: $\sum(\mathbf{x}\neq\mathbf{y})/n$ where $n$ is the bit length
  • Inner product similarity $\mathbf{x}\cdot\mathbf{y}$, particularly suitable for embeddings trained with dot-product-based loss functions
@@
For traditional binary hash codes (0s and 1s), we implement Hamming distance which calculates the number of differing bits between two hash codes efficiently on CPUs. The library also supports hash codes represented as -1s and 1s, enabling the use of inner product, cosine similarity, and Euclidean distance. This design choice allows users to leverage GPU acceleration for hash code comparisons, potentially boosting performance in GPU-centric environments. For continuous embeddings, all measures except Hamming distance are applicable.
@@
[…]
Recall is particularly suitable for instance-level retrieval as it measures whether specific, individual items have been found. For category-level retrieval, we use mean average precision at $k$ (mAP@k@k) as the standard evaluation metric. Precision at $k$ ($P@k$) is defined as follows:
[…]
For a single query, AP@k@k averages the precision values at all available ranks within the top $k$. For a query set, mAP@k@k computes the mean of the AP@k@k scores of all the queries within the query set. mAP@k@k is well-suited for evaluating category-level retrieval where queries typically have a large number of relevant items, as it measures a model’s ability to include many relevant results towards the top of a large retrieved set. Additionally, we implement the 11-point precision-recall curve for the category-level retrieval task, which plots interpolated precision at recall levels $R=\{0.0,0.1,\ldots,1.0\}$.